OpenStudy (anonymous):

Which measurement is equivalent to 4221 cm3? A. 4221 L B. 42.21 L C. 4.221 L D. 0.4221 L @Shay17

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@ganeshie8 pleez help I have one minute lest pleez

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@geerky42 plaaeeez pleez help

OpenStudy (anonymous):

@geerky42

OpenStudy (lyrae):

1 L = 1 dm^3 1 dm^3 = 1000 cm^3

OpenStudy (lyrae):

Which implies that: 1 L = 1000 cm^3 resulting in a conversion factor of: 1 cm^3 = 0.001 L

OpenStudy (anonymous):

that's not a choice bro @Lyrae

OpenStudy (lyrae):

Yes, I know. It's the relation (not conversion factor as I previously wrote) between cm^3 and liters (although you can use it to calculate the the conversion factor). Conversion factor: \[\frac{ 0.001 \space L }{ 1 \space cm^3 } = 0.001 \frac{L}{ cm^3 }\] To get the final answer: \[4221 \space cm^3 \times 0.001\frac{ L }{cm^3} = 4.221 \space L\]